OpenStudy (anonymous):

A manufacturer of bird feeders cuts spacers from a tube that is 3 1/3 inches long. How many spacers can be cut from the tube if each spacer must be 2/3 inch thick? estimated answer is ? exact answer is ?

ganeshie8 (ganeshie8):

# of spacers = \(\huge \frac{3\frac{1}{3}}{\frac{2}{3}}\)

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

so how would i simplify with fractions tho

ganeshie8 (ganeshie8):

maybe first simplify numerator : convert the "mixed fraction" into proper fraction...

OpenStudy (anonymous):

10/3/2/3 right

ganeshie8 (ganeshie8):

Correct ! simplify further

ganeshie8 (ganeshie8):

\(\huge \frac{3\frac{1}{3}}{\frac{2}{3}}\) \(\huge \frac{\frac{10}{3}}{\frac{2}{3}}\)

ganeshie8 (ganeshie8):

flip the denominator and write it as multiplication

ganeshie8 (ganeshie8):

\(\huge \frac{3\frac{1}{3}}{\frac{2}{3}}\) \(\huge \frac{\frac{10}{3}}{\frac{2}{3}}\) \(\huge \frac{10}{3} \times\frac{3}{2}\)

OpenStudy (anonymous):

5

ganeshie8 (ganeshie8):

5 is \(\large \color{red}{\checkmark}\)

ganeshie8 (ganeshie8):

good job !!

OpenStudy (anonymous):

is that th estmated answer of exact answer

ganeshie8 (ganeshie8):

5 is exact

OpenStudy (anonymous):

ok thanks

ganeshie8 (ganeshie8):

u wlc :) for estimated just do below : since we wanto roughly estimate quick, assume the tube is 3 inch long, then u wud get 3/2/3 = 9/2 = 4.5 = 4 spacers